Hello. Please help me ASAP- I need to turn this in quick. So I have a worksheet that says that I need to discriminate the solutions of the quadratic equation: x squared + 5x + 3 + a. How do I do it and what does the direction mean?
Thanks.

Answer by stanbon(75887)

discriminate the solutions of the quadratic equation: x squared + 5x + 3 + a.
-----------------
The discriminant of a quadratic is b^2-4ac
---------------
Your quadratic has a = 1, b=5, c=3+a
-----------
Your discriminant is 5^2 -4*1*(3+a) = 25-12-4a = 13-4a
---------------------
When the discriminant is less than zero you have no Real Number solutions:
13-4a < 0
4a>13
a > 13/4
-----------
When the discriminant is equal to zero you have two equal Real number solutions:
13-4a = 0
4a = 13
a = 13/4
-------------
when the discriminant is greater than zero you have two unequal Real number
solutions:
13-4a >0
4a<13
a < 13/4
===============
Cheers,
Stan H.